What is the wavelength of blue light that has a frequency of 6.37 X 10^14 s^1? (c= 3.00 X 10^8 m/s)
1
Expert's answer
2020-04-27T01:52:50-0400

Wavelength l is:

l = c/f = 3*10^8/6.37*10^14 = 4.71*10*-7 m = 471 nm